OpenStudy (anonymous):

A 100-cm3 container has 4 g of ideal gas in it at 250 kPa. If the volume is changed to 50 cm3 and the temperature remains constant, what is its new density?

OpenStudy (anonymous):

\[\rho=m/V=4g/50cm^3=0.08g/cm^3\]